A random variable follows the continuous uniform distribution between

20

and

50

​a) Calculate the probabilities below for the distribution.

​1)

​P(x≤40​)

​2)

​P(x=39​)

​b) What are the mean and standard deviation of this​ distribution?
Answer #1

X ~ U (20 , 50)

a) 1) P(X < 40) = (40 - 20) / (50 - 20) = 2/3 or 0.67

2) P(X = 39) = 0

b) mean = (20 + 50) / 2 = 35

standard deviation = (50 - 20) / sqrt(12) = 8.66